OpenStudy (anonymous):

Pre Calc: 29. For what value of k will the graph of 2x + ky = 6 be perpendicular to the graph of 6x – 4y = 12?

OpenStudy (anonymous):

recall that perpendicular implies that the slope M of the two lines you are comparing are "opposite reciprocals"

OpenStudy (anonymous):

So, start by getting the lines into the form y=mx+b y=-2x/k +6/k y=6x/4+12/-4 6/4=3/2 so the opposite reciprocal of this slope is -2/3

OpenStudy (anonymous):

k = 3.... Product of slopes should be -1(the lines are to be perpendicular) and slope of 6x-4y =12 is 3/2 and slope of perpendicular line is -2/3 which is possible iff k(coefficient of x) is 3....

OpenStudy (anonymous):

Thank you <3 <3